Solve 6x3 + 732 – 92 + 2 = 0
Answer #1

Ex3 +70²_ qut2=0. (x+2) (64²5x+1)= 0 (1+2)(n-1) (»–13) n=-2, 1/2,1/3 u= -2], (u= 1/2 In=1/3